1- For the following reaction, Na+(aq) + OH-(aq) + H+(aq) + Cl- (aq)  à Na+(aq) + Cl-...

1-

For the following reaction,

Na+(aq) + OH-(aq) + H+(aq) + Cl- (aq)  à Na+(aq) + Cl- (aq) + H2O (l)

Na+(aq) + Cl- (aq), ions not involved in the reaction, are called _________________ions.

2-

For the following reaction,

Na+(aq) + OH-(aq) + H+(aq) + Cl- (aq)  à Na+(aq) + Cl- (aq) + H2O (l)

What type of reaction is taking place? ____________________

3-

What is the mass in grams of 1.73 mol of MgBr2? __________

4-

What is the percentage composition of each element in PbCO3. Record each percentage to the tenths' place.

5-

What are the products for the reaction of hydrochloric acid and sodium sulfite? Answers MUST be written in the following order: metal salt + liquid + gas

6-

Which product will not be formed from a metathesis reaction?

Hydrobromic acid, HBr

Carbon Dioxide, CO2

Calcium Hydroxide, Ca(OH)2

Acetic acid, CH3COOH

7-

Find the percent yield of AgCl when a 5.95-g sample of AgNO3 is reacted with excess BaCl2 to give 4.48 g of AgCl, according to the equation:

2 AgNO3 (aq) + BaCl2 (aq) → 2AgCl (s) + Ba(NO3)2 (aq)

89.2 %

75.3 %

59.5 %

44.6 %
Answer #1

1) the ions which doesn't take place directly in the reaction are called spectator ions

2) this reaction is called neutralization reaction where an acid and a base reacts to form salt and water with release of energy.

3) molecular mass of MgBr2 = 184 g/mol

Mass= moles × molar mass

Mass = 1.73 × 184 = 318.32g

4) molar mass of PbCO3= 267g/mol

Mass of Pb= 207 g/mol

Percentage composition of Pb= (207/267)×100

Percentage composition of Pb= 77.5 %

Mass of C = 12

Percentage comp of C = (12/267)×100= 4.5

Percentage composition of oxygen =

Mass of 3 oxygen atoms = 3×16=48g/mol

Percentage comp = (48/267)×100= 18%